# ForAug: Recombining Foregrounds and Backgrounds to Improve Vision Transformer Training with Bias Mitigation

### v1

@@ 4.1 Design Choices of ForAug
Note that $s\circ s=\mathrm{id}$ on $[0,1]$. This way, distributions with $\eta\leq-1$ are more concentrated around the borders. $\eta=1$ and $\eta=-1$ both correspond to the uniform distribution. The PDF of this extended Bates distribution is visualized in Figure 3.
When sampling more towards the center of the image, the difficulty of the task is reduced, which then reduces the performance on TinyImageNet. This is reflected in the performance when evaluating on TinyForNet with $\eta=2$ and $\eta=3$ compared to $\eta=-1/1$. We observe a similar reduction for $\eta<-1$. This experiment is conducted using the LaMa infill model.
Dataset Classes Training Images Validation Images
TinyImageNet 200 100,000 10,000
TinyForNet 200 99,404 9,915
ImageNet 1,000 1,281,167 50,000
ForNet 1,000 1,274,557 49,751
Table 3: Dataset statistics for TinyImageNet, TinyForNet, ImageNet, and ForNet. For ForNet and TinyForNet we report the number of foreground/background pairs.
After fixing the optimal design parameters in Table 1 (last row), we construct the full ForNet dataset using the entire ImageNet dataset. Table 3 compares the dataset statistics of ImageNet and ForNet. The slightly reduced image count in ForNet is due to instances where Grounded SAM failed to produce valid object detections.
4.2 Image Classification Results
Model ImageNet Accuracy when trained on Delta
ImageNet ForNet
ViT-S $79.1\pm 0.1$ $81.4\pm 0.1$ +2.3
ViT-B $77.6\pm 0.2$ $81.1\pm 0.4$ +3.5
ViT-L $75.3\pm 0.4$ $79.8\pm 0.1$ +4.5
Swin-Ti $77.9\pm 0.2$ $79.7\pm 0.1$ +1.8
Swin-S $79.4\pm 0.1$ $80.6\pm 0.1$ +1.2
ResNet-50 $78.3\pm 0.1$ $78.8\pm 0.1$ +0.5
ResNet-101 $79.4\pm 0.1$ $80.4\pm 0.1$ +1.0
Table 4: ImageNet results of models trained on ForNet and on ImageNet directly. ForNet improves the performance of all models in our test.
Table 4 compares the ImageNet performance of models trained on ForNet and ones trained directly on ImageNet. We adopt the training setup of [33] and [48] (details in Appendix A) for training ViT [10], Swin [31] and ResNet [16] models. Notably, ForNet improves performance across all tested architectures, including the ResNet models (up to $1$ p.p.), demonstrating benefits beyond Transformers. For Transformer models, we observe improvements from $1.2$ p.p. to $4.5$ p.p. This improvement is more substantial for the larger models, with ViT-L gaining $4.5$ p.p. in accuracy. ForNet’s improvements mostly counteract the drop in performance due to overfitting for large models. When training on ImageNet, this drop is $3.8$ p.p. from ViT-S to ViT-L, while for ForNet it is reduced to $1.6$ p.p.
Model Aircraft Cars Flowers Food Pets
ViT-S @ ImageNet $72.4\pm 1.0$ $89.8\pm 0.3$ $94.5\pm 0.2$ $89.1\pm 0.1$ $93.8\pm 0.2$
ViT-S @ ForNet $78.6\pm 0.5$ $92.2\pm 0.2$ $95.5\pm 0.2$ $89.6\pm 0.1$ $94.5\pm 0.2$
+6.2 +2.4 +1.0 +0.5 +0.7
ViT-B @ ImageNet $71.7\pm 0.5$ $90.0\pm 0.2$ $94.8\pm 0.4$ $89.8\pm 0.2$ $94.1\pm 0.4$
ViT-B @ ForNet $79.0\pm 2.2$ $93.3\pm 0.1$ $96.5\pm 0.1$ $90.9\pm 0.1$ $95.1\pm 0.4$
+7.3 +3.3 +1.7 +1.1 +1.0
ViT-L @ ImageNet $72.1\pm 1.0$ $88.8\pm 0.3$ $94.4\pm 0.3$ $90.1\pm 0.2$ $94.2\pm 0.4$
ViT-L @ ForNet $77.6\pm 1.2$ $89.1\pm 0.2$ $96.6\pm 0.1$ $91.3\pm 0.1$ $95.1\pm 0.1$
+5.5 +0.3 +2.2 +1.2 +0.9
Swin-Ti @ ImageNet $77.0\pm 0.1$ $91.3\pm 0.6$ $95.9\pm 0.1$ $90.0\pm 0.2$ $94.2\pm 0.1$
Swin-Ti @ ForNet $81.1\pm 0.8$ $92.8\pm 0.4$ $96.2\pm 0.1$ $90.4\pm 0.3$ $94.8\pm 0.5$
+4.1 +2.5 +0.3 +0.4 +0.6
Swin-S @ ImageNet $75.7\pm 1.4$ $91.0\pm 0.3$ $95.9\pm 0.5$ $91.1\pm 0.2$ $94.4\pm 0.1$
Swin-S @ ForNet $81.4\pm 0.2$ $93.1\pm 0.2$ $96.3\pm 0.3$ $91.2\pm 0.2$ $94.9\pm 0.3$
+5.7 +2.1 +1.4 +0.1 +0.5
ResNet-50 @ ImageNet $78.2\pm 0.5$ $89.8\pm 0.2$ $91.7\pm 0.4$ $84.4\pm 0.2$ $93.7\pm 0.3$
ResNet-50 @ ForNet $80.3\pm 0.4$ $90.4\pm 0.2$ $91.7\pm 0.2$ $84.5\pm 0.2$ $93.7\pm 0.3$
+2.1 +0.6 $\pm$0 +0.1 $\pm$0
ResNet-101 @ ImageNet $78.4\pm 0.6$ $90.3\pm 0.1$ $91.2\pm 0.5$ $86.0\pm 0.2$ $94.3\pm 0.2$
ResNet-101 @ ForNet $81.4\pm 0.5$ $91.3\pm 0.1$ $92.9\pm 0.2$ $86.3\pm 0.1$ $94.0\pm 0.3$
+3.0 +1.3 +1.7 +0.3 -0.3
Table 5: Downstream accuracy in percent when finetuning on other datasets. Models were pretrained on ForNet and ImageNet. Pretraining on ForNet increases Transformer downstream accuracy on all datasets.
To assess the transferability of ForNet-trained models, we finetune models pretrained on ImageNet and ForNet on five fine-grained datasets: FGVC-Aircraft [32], Stanford Cars [7], Oxford Flowers [34], Food-101 [21], and Oxford-IIIT Pets [35]. While for ResNets, the performance of both training datasets is about the same, for every Transformer, we see the accuracy improve on all downstream dataset by up to 7.3 p.p. and a reduction of error rate of up to $39.3\%$. In summary, these results demonstrate that the improved representation learning achieved by training on ForNet translates to superior performance not only on ImageNet, but also on a variety of fine-grained image classification tasks.
4.3 Further Model Evaluation
Beyond its use for training, ForNet’s unique properties and controlled data generation capabilities make it a powerful tool for analyzing model behavior and biases.
Background Robustness
Model Background Robustness when trained on Delta
ImageNet ForNet
ViT-S $0.73\pm 0.01$ $0.99\pm 0.01$ +0.26
ViT-B $0.72\pm 0.01$ $1.00\pm 0.01$ +0.28
ViT-L $0.70\pm 0.01$ $1.00\pm 0.01$ +0.30
Swin-Ti $0.72\pm 0.01$ $1.00\pm 0.01$ +0.28
Swin-S $0.72\pm 0.01$ $1.00\pm 0.01$ +0.28
ResNet-50 $0.79\pm 0.01$ $0.99\pm 0.01$ +0.20
ResNet-101 $0.79\pm 0.01$ $1.00\pm 0.01$ +0.21
Table 6: Evaluation of the background robustness of models trained on ForNet and on ImageNet directly. Training on ForNet improves the background robustness of all model to $\approx 1.00$, meaning the model is indifferent to the choice of background.
We assess the robustness of models to shifts in the background distribution from a class-related background to any background. Background robustness is defined to be the ratio of accuracy on ForNet with same-class backgrounds to accuracy with any background:
Background $\text{Background Robustness}=\frac{\text{Acc}(\textit{ForNet}_{\text{all}})}{\text{Acc}(\textit{ForNet}_{\text{same}})}$ (4)
It represents the relative drop in performance under a background distribution shift. Table 6 presents the background robustness of various models. When trained on ImageNet, smaller models generally exhibit greater robustness to changes in the background distribution than larger models and ResNet is more robust than the tested Transformer models. Crucially, training on ForNet instead of ImageNet improves the background robustness of all models to $\approx 1.00$, meaning that these models are agnostic to the choice of background and only classify based on the foreground. These findings highlight the generalization benefits of ForNet.
Foreground Focus
Model Foreground Focus when trained on
IN FN IN FN IN FN
GradCam GradCam++ IG
ViT-S $1.2\pm 0.1$ $2.3\pm 0.3$ $1.2\pm 0.1$ $2.1\pm 0.4$ $1.9\pm 0.1$ $2.7\pm 0.1$
ViT-B $1.2\pm 0.1$ $2.4\pm 0.7$ $1.1\pm 0.1$ $2.1\pm 0.1$ $1.7\pm 0.1$ $2.7\pm 0.1$
ViT-L $1.3\pm 0.1$ $1.6\pm 0.1$ $1.1\pm 0.1$ $1.3\pm 0.1$ $1.3\pm 0.1$ $2.6\pm 0.1$
Swin-Ti $0.9\pm 0.1$ $0.7\pm 0.1$ $1.0\pm 0.3$ $0.7\pm 0.3$ $2.5\pm 01$ $4.8\pm 0.3$
Swin-S $0.8\pm 0.1$ $0.7\pm 0.1$ $0.7\pm 0.1$ $0.7\pm 0.4$ $2.4\pm 0.1$ $4.6\pm 0.3$
ResNet-50 $2.2\pm 0.1$ $2.7\pm 0.1$ $2.0\pm 0.1$ $2.9\pm 0.1$ $3.2\pm 0.1$ $4.9\pm 0.2$
ResNet-101 $2.3\pm 0.1$ $2.8\pm 0.1$ $2.2\pm 0.1$ $3.0\pm 0.1$ $3.2\pm 0.1$ $4.8\pm 0.1$
Table 7: Evaluation of the foreground focus using GradCam, GradCam++ and IntegratedGradients of models trained on ForNet (FN) and on ImageNet (IN) directly. Training on ForNet improves the foreground focus of almost all models.
Leveraging our inherent knowledge of the foreground masks when using ForNet, as well as common XAI techniques [40, 4, 44], we can evaluate a model’s focus on the foreground object. We can directly evaluate ImageNet trained models, but this technique can also be extended to other datasets without relying on manually annotated foreground-masks. To evaluate the foreground focus, we employ Grad-CAM [40], Grad-CAM++ [4] or IntegratedGradients (IG) [44] to compute the per-pixel importance of an image for the model’s prediction. The foreground focus is defined to be the ratio of the foreground’s relative importance to its relative size in the image:
[…]
The foreground focus of a model is its average foreground focus over all test images. Table 7 presents our findings. Training on ForNet significantly increasees the foreground focus of ViT and ResNet across all metrics used. For Swin, the foreground focus stagnates when measured using GradCam and GradCam++, but almost doubles when using IG.
Center Bias
Model Center Bias when trained on Delta
ImageNet ForNet
$0.255\pm 0.008$ $0.220\pm 003$0.035
$0.254\pm 0.004$ $0.190\pm 0.002$0.064
$0.243\pm 0.011$ $0.117\pm 0.007$0.126
$0.250\pm 0.007$ $0.165\pm 0.002$0.085
$0.232\pm 0.001$ $0.156\pm 002$0.076
$0.263\pm 0.003$ $0.197\pm 0.003$0.066
$0.230\pm 0.003$ $0.199\pm 002$0.031
Table 8: Evaluation of the position bias. We plot the accuracy relative to the center accuracy of multiple instantiations of the models when the foreground objects is in different cells a $3\times 3$ grid. Training on ForNet significantly reduces a models center bias.
With ForNet we have unique control over the position of the foreground object in the image. This lets us quantify the center bias of ImageNet- and ForNet-trained models. We divide the image into a $3\times 3$ grid and evaluate model accuracy when the foreground object is in each of the $9$ grid cells. Each cell’s accuracy is divided by the accuracy in the center cell for normalization, which gives us the relative performance drop when the foreground is in each part of the image. The center bias is calculated as one minus the average of the minimum performance of a corner cell and the minimum performance of a side cell:
Center $\begin{split}&\text{Center Bias}=\\ &\hskip 7.0pt1-\frac{\min\limits_{a,b\in\{0,2\}}\text{Acc}(\text{cell}_{(a,b)})+\min\limits_{\begin{subarray}{c}a=1\text{ or }b=1\\ a\neq b\end{subarray}}\text{Acc}(\text{cell}_{(a,b)})}{2\text{Acc}(\text{cell}_{(1,1)})}\end{split}$ (6)
Table 8 visualizes the center bias of three instantiations of each model. Performance is generally highest in the center and the center top and bottom and center left and right cells, and lowest in the four corners. Interestingly, ImageNet-trained models perform slightly better when the foreground object is on the right side of the image, compared to the left side, despite our use of random flipping with a probability of $0.5$ during training. Training on ForNet significantly reduces center bias across all models. This demonstrates that ForNet promotes a more uniform spatial attention distribution. Their accuracy is higher in the center left and right cells than in the center top and bottom ones, which is not the case for ImageNet-trained models.
Size Bias
Figure 4: Evaluation of the size bias of models trained on ForNet. We plot the accuracy relative to the accuracy when using the mean foreground size.
Finally, we evaluate the impact of different-sized foreground objects on the accuracy. For this evaluation, we use the mean foreground size strategy. We introduce a size factor $f_{\text{size}}$ by which we additionally scale the foreground object before pasting it onto the background. Results are again normalized by the accuracy when using the mean foreground size ($f_{\text{size}}=1.0$). Figure 4 shows the size bias curves of ViT-S and ViT-B when trained on ImageNet and ForNet. Models trained on ForNet maintain better performance even with smaller foreground objects, when ImageNet-trained models exhibit a more rapid performance decline. Therefore, ForNet-training improves robustness to variations in object scale.
5 Discussion & Conclusion
We introduce ForAug, a novel data augmentation scheme that facilitates improved Transformer training for image classification. By explicitly separating and recombining foreground objects and backgrounds, ForAug enables controlled data augmentation, leading to significant performance gains on ImageNet and downstream fine-grained classification tasks. Furthermore, ForAug provides a powerful framework for analyzing model behavior and quantifying biases, including background robustness, foreground focus, center bias, and size bias. Our experiments demonstrate that training on ForNet, the instantiation of ForAug on ImageNet, not only boosts accuracy but also significantly reduces these biases, resulting in more robust and generalizable models. In the future, we see ForAug be also applied to other datasets and tasks, like video recognition or segmentation. ForAug’s ability to both improve performance and provide insights into model behavior makes it a valuable tool for advancing CV research and developing more reliable AI systems.
Acknowledgements
This work was funded by the Carl-Zeiss Foundation under the Sustainable Embedded AI project (P2021-02-009) and by the EU project SustainML (Horizon Europe grant agreement No 101070408). All compute was done thanks to the Pegasus cluster at DFKI.
Appendix A Training Setup
@@
Image Resolution $224\times 224$
Epochs 300
Learning Rate 3e-3
Learning Rate Schedule cosine decay
Batch Size 2048
Warmup Schedule linear
Warmup Epochs 3
Weight Decay 0.02
Label Smoothing 0.1
Optimizer Lamb [57]
Data Augmentation Policy 3-Augment [48]
Table 9: Training setup for our ImageNet and ForNet training.
Dataset Batch Size Epochs Learning Rate
Aircraft 512 500 3e-4
Cars 1024 500 3e-4
Flowers 256 500 3e-4
Food 2048 100 3e-4
Pets 512 500 3e-4
Table 10: Training setup for finetuning on different downstream datasets. Other settings are the same as in Table 9.
On ImageNet we use the same training setup as [33] and [48] without pretraining. As our focus is on evaluating the changes in accuracy due to ForAug/ForNet, like [33], we stick to one set of hyperparameters for all models. We list the settings used for training on ImageNet and ForNet in Table 9 and the ones used for finetuning those weights on the downstream datasets in Table 10.
Appendix B Infill Model Comparison
@@
Table 11: Example infills of LaMa and Attentive Eraser.
Appendix C Images with High Infill Ratio

### v2

@@ Appendix A Extended Bates Distribution
This function effectively maps the central portion of the interval to the edges and the edge portions to the center. For example, a value of 0.3 (central-left) is mapped to 0.8 (edge-right), while 0.8 (edge-right) is mapped to 0.3 (central-left). This transformation inverts the distribution’s concentration, shifting the probability mass from the center to the borders. We visualize the distribution function of the extended Bates distribution in Figure 6. Both $\eta=1$ and $\eta=-1$ result in a uniform distribution across the image.
Appendix B Resource Usage of ForAug
To utilize the proposed ForAug, specific computational resources are necessary, particularly for computing and storing for the output of the segmentation stage and for on-the-fly processing of the recombination stage.
Segmentation.
ForAug involves a computationally expensive segmentation and infill stage, which is a one-time calculation per dataset. Once computed, the segmentation and infill results can be perpetually reused, amortizing the initial cost over all subsequent experiments and applications. On NVIDIA H100 GPUs, the segmentation stage will compute at a rate of $374.3\frac{\text{img}}{\text{GPU}\times\text{h}}$ when using Attentive Eraser or $5338.6\frac{\text{img}}{\text{GPU}\times\text{h}}$ for LaMa. For ImageNet this comes down to just under 9 days (Attentive Eraser) or 16 hours (LaMa) on two 8 GPU nodes. To facilitate immediate use and reproduction of results, we publicly provide the precalculated segmentation stage output for the ImageNet dataset for download [^2] Link will go here.. The output of ForAug’s segmentation step on ImageNet dataset requires 73 GB of additional disk space for the segmentation output, which is separate from the base 147 GB ImageNet size.
Recombination.
The recombination step of ForAug is implemented as a based data loader operation. It’s thus offloaded to the CPU, where it can be heavily parallelized and thus only results in a very minor increase in the training step-time. For example, using a ViT-B model on an NVIDIA A100 GPU, the average update step-time increased by $1\%$, from $528\pm 2$ ms to $534\pm 1$ ms.
Appendix C Training Setup
Table 9: Training setup and hyperparameters for our ImageNet training.
Parameter ViT, Swin, ResNet DeiT
Image Resolution $\times 224 224×224$$\times 224$
Epochs 300 300
Learning Rate 3e-3 S/B: 1e-3, L: 5e-4
Learning Rate Schedule cosine decay cosine decay
Batch Size 2048 1024
GPUs $4\times$ NVIDIA A100/H100/H200 $4\times$ NVIDIA A100/H100/H200
Warmup Schedule linear linear
Warmup Epochs 3 3
Weight Decay 0.02 0.05
Label Smoothing 0.1 0.1
Optimizer Lamb [59] AdamW
Data Augmentation Policy 3-Augment [48] DeiT [47]
Augmentations
Resize
RandomCrop
HorizontalFlip
Grayscale
Solarize
GaussianBlur
ColorJitter
CutMix [61]
RandomResizedCrop
HorizontalFlip
RandomErase [64]
RandAugment [5]
ColorJitter
Mixup [63]
CutMix [61]
Table 10: Training setup for finetuning on different downstream datasets. Other settings are the same as in Table 9. For finetuning, we always utilize 3-Augment and the related parameters from the ViT, Swin, ResNet column of Table 9
Dataset Batch Size Epochs Learning Rate Num. GPUs
Aircraft 512 500 3e-4 2
Cars 1024 500 3e-4 4
Flowers 256 500 3e-4 1
Food 2048 100 3e-4 4
Pets 512 500 3e-4 2
On ImageNet we use the same training setup as [32] and [48] without pretraining for ViT, Swin, and ResNet. For DeiT, we train the same ViT architecture but using the data augmentation scheme and hyperparameters from [47]. As our focus is on evaluating the changes in accuracy due to ForAug, like [32], we stick to one set of hyperparameters for all models. We list the settings used for training on ImageNet in Table 9 and the ones used for finetuning those weights on the downstream datasets in Table 10. Out implementation is using PyTorch [35] and the timm library [55] for model architectures and basic functions.
Table 11: Hardware and Software specifics used for both training and evaluation.
@@ Appendix C Training Setup
GPU NVIDIA A100/H100/H200
CPU 24 CPU cores (Intex Xenon) per GPU
Memory up to 120GB per GPU
Operating System Enroot container for SLURM based on Ubuntu 24.04 LTS
Python 3.12.3
PyTorch 2.7.0
TorchVision 0.22.0
Timm 1.0.15
Table 11 lists the specific hardware we use, as well as versions of the relevant software packages.
Appendix D Infill Model Comparison
Table 12: Example infills of LaMa and Attentive Eraser.
@@ Appendix D Infill Model Comparison
We visualize example infilled images for both LaMa [45] and Attentive Eraser [43] in Table 12. We qualitatively find that while LaMa often leaves repeated textures of blurry spots where the object was erased, Attentive Eraser produces slightly cleaner and more coherent infills of the background.
Appendix E Image Infill Ratio
@@
Figure 7: We plot the distribution of the relative size of the detected foreground object that is infilled in our Segmentation step of ImageNet. While most images contain objects of smaller size, there is a peak where Grounded SAM [37] detects almost the whole image as the foreground object. For examples of such large infills, see Table 13.
Table 13 shows infills for images where Grounded SAM [37] marks a high percentile of the image as the foreground object (Infill Ratio), that has to be erased by the infill models. While LaMa tends to fill those spots with mostly black or gray and textures similar to what we saw in Table 12, Attentive Eraser tends to create novel patterns by copying what is left of the background all over the rest of the image. Figure 7 plots the distribution of infill ratios in ForAug. While there is a smooth curve of the number of detections decreasing with the infill ratio until $\approx 90\%$, there is an additional peak at $\approx 100\%$ infill ratio. We believe that this peak is made up of failure cases of Grounded SAM.
We filter out all backgrounds that have an infill ratio larger than our pruning threshold $t_{\text{prune}}=0.8$, which translates to $10\%$ of backgrounds.